OpenStudy (anonymous):

At a local grocery, 26-ounces of a sport drink costs $1.30. What is the price per ounce? A) $0.03 B) $0.05 C) $0.30 D) $2.04

OpenStudy (anonymous):

1.30/26= .05

OpenStudy (anonymous):

Paige has to make invitation cards for a party. If she can make 18 cards in 2 days, how many cards can she make in 30 days? A) 540 B) 300 C) 270 D) 30

OpenStudy (anonymous):

30/2=15 15*18=270 B) 270

OpenStudy (anonymous):

Cheryl drove 4 hours at a constant rate to travel 160 miles. At what rate did she drive? A) 20 mph B) 35 mph C) 40 mph D) 72 mph

OpenStudy (anonymous):

160/4= 40

OpenStudy (anonymous):

Ralph runs at a constant rate of 3 miles per hour. How long will it take Ralph to run 15 miles? A) 2 hr. B) 3 hr. C) 4 hr. D) 5 hr.

OpenStudy (anonymous):

15/3= 5 D) 5 hr.

OpenStudy (anonymous):

what grade and math level are you in?

OpenStudy (anonymous):

i didnt on my own an i got c grrr i dont get math 7th grade i got intensive math and advance wich makes no sense to me

OpenStudy (anonymous):

this is realllly easy stuff. like 3rd grade stuff...